### v2

@@ Abstract
Video localization tasks aim to temporally locate specific instances in videos, including temporal action localization (TAL), sound event detection (SED) and audio-visual event localization (AVEL). Existing methods over-specialize on each task, overlooking the fact that these instances often occur in the same video to form the complete video content. In this work, we present UniAV, a Unified Audio-Visual perception network, to achieve joint learning of TAL, SED and AVEL tasks for the first time. UniAV can leverage diverse data available in task-specific datasets, allowing the model to learn and share mutually beneficial knowledge across tasks and modalities. To tackle the challenges posed by substantial variations in datasets (size/domain/duration) and distinct task characteristics, we propose to uniformly encode visual and audio modalities of all videos to derive generic representations, while also designing task-specific experts to capture unique knowledge for each task. Besides, we develop a unified language-aware classifier by utilizing a pre-trained text encoder, enabling the model to flexibly detect various types of instances and previously unseen ones by simply changing prompts during inference. UniAV outperforms its single-task counterparts by a large margin with fewer parameters, achieving on-par or superior performances compared to state-of-the-art task-specific methods across ActivityNet 1.3, DESED and UnAV-100 benchmarks.
Keywords:
Multi-modal video localization Multi-task learning
1 Introduction
@@
With the explosion of video content due to social networks and digital cameras, video understanding [27, 59, 12, 50] continues to be one of the essential research domains in computer vision. Videos recorded in natural scenes are always untrimmed and comprise both visual and audio modalities. They usually cover multiple instances of interest, including visible actions, audible sound events as well as audio-visual events [47] that are both audible and visible at the same time. For example, as illustrated in Fig. 1, we can discern the visual action of “playing ten pins”, the audio-visual event of “striking bowling”, and also the background narration of “man/woman speaking”. All these events are equally crucial, jointly contributing to the overall understanding of video content.
@@
Figure 1: Our unified framework can localize all three kinds of instances in untrimmed videos, including visual actions, sound events and audio-visual events. All these instances contribute to a comprehensive understanding of video content.
@@
However, current video localization approaches only concentrate on recognizing and detecting one type of these instances, involving the tasks of temporal action localization (TAL) [26, 59, 30], sound event detection (SED) [20, 52] and audio-visual event localization [47, 50, 11] (AVEL). Despite convenience for some specific applications, such separate definitions bring the following drawbacks: 1) Independent designs cause redundant parameters since recent localization models usually adopt similar architectures, e.g., transformer backbones in [59, 30, 11]. 2) It hinders models from learning and sharing generic knowledge between different tasks and modalities. For example, rich TAL data enables models to identify common instances, which can naturally assist AVEL and SED tasks. Additionally, AVEL data allows models to learn corresponding audio representations for many visual actions in TAL and learn visual cues for sound events in SED, thereby facilitating improvements in both TAL and SED tasks. Besides, some recent task-specific methods [1, 21, 15] have verified that integrating visual and audio modalities is beneficial for TAL and SED tasks.
@@
In this work, we aim to develop a unified framework to localize all these types of instances in untrimmed videos, solving three video localization tasks (TAL, AVEL and SED) by a single model. However, the main obstacles hindering this attempt lie in two aspects. Firstly, the datasets for these tasks exhibit distinct properties with significant domain and duration gaps. For example, ActivityNet [6] for TAL focuses on human activities, while UnAV-100 [11] for AVEL and DESED [48] for SED contain events from other domains such as animals, nature and tools. Besides, the instance duration of different datasets varies greatly, e.g., over 50% events in DESED are less than 1s, while the instances in ActivityNet and UnAV-100 are much longer, with the longest ones lasting over 7 minutes. Secondly, different tasks emphasize different video characteristics and modalities. TAL pays more attention to capturing temporal relationships of actions on the visual track, while SED is dedicated to the fine-grained understanding of sound events. In contrast, AVEL assigns equal importance to both auditory and visual cues. Thus, unifying these three tasks is challenging, and just simple joint training may lead to a significant decrease in performance.
@@
To tackle these challenges, we introduce UniAV, a multi-task learning framework for Unified Audio-Visual perception in video localization tasks. We unify TAL, SED and AVEL tasks within a single model from three aspects. 1) Unified audio-visual encoding. In order to unify diversity between the data from different tasks and obtain general input representations, we employ the large pre-trained model [49] to uniformly tokenize the visual and audio modalities of all input videos. Then, the obtained embeddings are fed into an audio-visual pyramid transformer network, enabling the model to detect both very short as well as long instances that span minutes. 2) Task-specific experts. Due to the divergence of different tasks, learning distinct knowledge for each task is critical. Thus, we design task-specific expert layers in our transformer blocks to learn task-specific features by switching to corresponding experts according to the input data. 3) Unified language-aware classifier. Datasets for different tasks pose their own category sets. Instead of using separate task-specific classification heads, we propose a unified language-aware classifier by tokenizing the class vocabularies with task-specific prompts using the pre-trained text encoder. Benefiting from this new formulation, our model gains the flexibility to detect different types of instances by simply changing prompts and expands the capability to recognize previously unseen instances.
@@
With the unified framework, UniAV can learn from diverse task-specific data and handle three video localization tasks with the same model parameters. Extensive experiments demonstrate that UniAV outperforms its single-task counterparts by a large margin with fewer parameters. Besides, multi-task joint training can be an effective pretraining step for the single-task models, leading to further gains and setting new state-of-the-art results across all three tasks, i.e., ActivityNet 1.3 [6] ($36.2\%$ average mAP) for TAL, DESED [20] ($61.1\%$ average mAP) for SED, and UnAV-100 [11] ($51.7\%$ average mAP) for AVEL.
@@
  • To the best of our knowledge, our UniAV is the first unified framework that solves temporal action localization, sound event detection and audio-visual event localization within a single model, leading to a holistic understanding of video content in real-world scenarios.
@@
  • We propose a unified audio-visual encoding pipeline to address data discrepancies across diverse tasks, while also incorporating task-specific experts to capture distinct features for each task.
@@
  • We design a unified language-aware classifier, allowing the model to flexibly detect various types of instances and previously unseen ones during inference.
2 Related Work
2.1 Temporal Localization Tasks
@@
Temporal action localization (TAL) aims to localize and classify action instances occurring in an untrimmed video. Supervised learning-based TAL can be categorized into two-stage [26, 27, 2, 55] and single-stage [24, 30, 45, 59] methods. However, previous works mainly focus on temporal modeling within the visual modality (e.g., RGB and optical flow), ignoring the information in its corresponding audio track. Recently, some works [1, 21, 22] have attempted to utilize the audio modality in videos for TAL, and have found it very helpful to detect the actions with strong audio cues, thus boosting the model performance.
@@
Sound event detection (SED) is a popular task in the audio signal processing community, which involves temporally detecting sound events in a purely acoustic scene. The DCASE Challenge [34] examines sound event detection in domestic environments as one of the challenge tasks [19, 20, 52]. Sound events typically come with their corresponding visual information, e.g., man speaking, playing guitar, and dog barking. It has been verified that incorporating visual modality is beneficial for SED in some recent works [15, 5].
@@
Audio-visual event localization (AVEL) aims to detect events that are simultaneously audible and visible in video content. Tian et al. [47] introduced the first AVE dataset and proposed an audio-guided visual attention model for the task. Subsequent studies [50, 54, 56, 10] primarily concentrated on event information modeling and cross-modal fusion strategies. However, these works formulate AVEL as a segment-level classification problem based on trimmed, short video clips. Each video clip only contains a single audio-visual event, which deviates from real-life scenarios involving diverse untrimmed videos. To solve the problem, Geng et al. built the UnAV-100 [11] dataset for localizing audio-visual events in untrimmed videos and proposed a model to recognize multiple events and regress their temporal boundaries in a single pass. Furthermore, few approaches for audio-visual video parsing [46, 29, 36, 38] and multisensory temporal event localization [17] strive to identify audio-only, visual-only and audio-visual events in videos. However, they are all segment-level classification methods with fragmented task definitions, and confined to weakly-supervised settings due to the lack of temporal annotations in videos during training. In comparison, we develop a multi-task supervised framework that learns multi-modal event localization from large-scale individually collected datasets for TAL, SED and AVEL with rich label vocabularies. Moreover, our model is proposal-based, capable of flexibly regressing and recognizing all temporal instances in untrimmed videos.
2.2 Multi-Task Learning
@@
Instead of separate training for each individual task, multi-task learning involves tackling multiple related tasks simultaneously, aiming to share and leverage knowledge across them. For example, recent works [33, 23, 44] proposed to jointly learn visual grounding tasks in a collaborative model. Yan et al. [57] presented Unicorn to solve four object tracking problems. Lu et al. [32] undertook training across 12 vision-language tasks, with each task having its own task-specific prediction head. For video understanding tasks, UniVTG [25] stands out for its emphasis on unifying three temporal grounding tasks: moment retrieval, highlight detection, and video summarization, employing specific query types. Addressing episodic memory tasks, MINOTAUR [13] exhibits proficiency in handling three egocentric vision tasks [14] through a singular model. In this work, we aim to temporally localize diverse modality-aware instances in untrimmed videos, thereby advancing comprehensive video understanding. This work pioneers the integration of temporal action localization, sound event detection, and audio-visual event localization within a unified framework for the first time.
3 The UniAV Framework
@@
Our goal is to develop a unified framework to localize visual actions, sound events and audio-visual events in an untrimmed video. To achieve this, we propose to unite three video localization tasks: TAL, SED and AVEL, leveraging inherent similarities among them. Figure 2 shows an overview of the proposed framework.
@@
Figure 2: The overview of our unified framework. Given a visual and audio pair from an untrimmed video, we first tokenize them by a pair of visual and audio encoders. Then, the encoded features are fed into an audio-visual pyramid transformer for cross-modal fusion at multiple temporal scales. The task-specific experts in transformer blocks learn distinct knowledge for each task. And the categories of each task are encoded with prompts to compute similarities with pyramid features, which are used to perform language-aware classification. Finally, the model recognizes classes and regresses temporal boundaries for all types of instances occurring in the video.
@@
Problem setting. Given an untrimmed video containing visual and audio tracks, our model aims to output the categories as well as the start/end timestamps of all instances occurring in the video. We formulate all three tasks (TAL, SED and AVEL) as a joint classification and regression problem. Formally, we first divide the input video into $T$ visual snippets $\{V_{t}\}_{t=1}^{T}$ and audio snippets $\{A_{t}\}_{t=1}^{T}$, where $T$ varies across videos. Following the setting [59, 11], the model is expected to predict the event label $Y=\{y_{i}\}_{i=1}^{N}$, where $N$ is the number of predicted events. Each instance $y_{i}=(s_{i},e_{i},c_{i})$ is defined by its onset $s_{i}$, offset $e_{i}$ and category $c_{i}$, where $s_{i},e_{i}\in[1,T]$, $c_{i}\in\{1,\dots,C\}$ is the pre-defined category.
3.1 Unified Audio-Visual Encoding

### v3

@@ Abstract
Video event localization tasks include temporal action localization (TAL), sound event detection (SED) and audio-visual event localization (AVEL). Existing methods tend to over-specialize on individual tasks, neglecting the equal importance of these different events for a complete understanding of video content. In this work, we aim to develop a unified framework to solve TAL, SED and AVEL tasks together to facilitate holistic video understanding. However, it is challenging since different tasks emphasize distinct event characteristics and there are substantial disparities in existing task-specific datasets (size/domain/duration). It leads to unsatisfactory results when applying a naive multi-task strategy. To tackle the problem, we introduce UniAV, a Unified Audio-Visual perception network to effectively learn and share mutually beneficial knowledge across tasks and modalities. Concretely, we propose a unified audio-visual encoder to derive generic representations from multiple temporal scales for videos from all tasks. Meanwhile, task-specific experts are designed to capture the unique knowledge specific to each task. Besides, instead of using separate prediction heads, we develop a novel unified language-aware classifier by utilizing semantic-aligned task prompts, enabling our model to flexibly localize various instances across tasks with an impressive open-set ability to localize novel categories. Extensive experiments demonstrate that UniAV, with its unified architecture, significantly outperforms both single-task models and the naive multi-task baseline across all three tasks. It achieves superior or on-par performances compared to the state-of-the-art task-specific methods on ActivityNet 1.3, DESED and UnAV-100 benchmarks.
I Introduction
@@
With the explosion of video content due to social networks and digital cameras, video understanding [1, 2, 3, 4, 5] continues to be one of the essential research domains in computer vision. Videos recorded in natural scenes are always untrimmed and comprise both visual and audio modalities. They usually cover multiple instances of interest, including visible-only actions, audible-only sound events as well as audio-visual events [6] that are both audible and visible at the same time. Localizing these various instances automatically using audio and visual signals in a video has wide practical applications including intelligent surveillance, human behavior analysis and human-robot interaction. As illustrated in Fig. 1, we can discern the visual action of “playing ten pins”, the audio-visual events of “striking bowling”, “man speaking”, and also the out-of-screen narration of “man/woman speaking”. All these events are equally crucial, jointly contributing to the overall understanding of video content.
@@
Figure 1: We propose a unified audio-visual perception framework to localize all three kinds of instances in untrimmed videos, including visual actions, sound events and audio-visual events. All these instances equally contribute to the comprehensive understanding of video content.
@@
However, existing video localization approaches only concentrate on recognizing and detecting one type of these instances, involving the tasks of temporal action localization (TAL) [7, 5, 8, 9, 10], sound event detection (SED) [11, 12] and audio-visual event localization (AVEL) [6, 13, 4, 14, 15]. Despite convenience for some specific applications, such separate definitions bring the following drawbacks: 1) Independent designs cause redundant parameters since recent localization models usually adopt similar architectures, e.g., transformer backbones in [8, 9, 15]. 2) It hinders models from learning and sharing mutually beneficial knowledge between different tasks and modalities. For example, rich TAL data enables models to identify common instances, which can naturally assist AVEL and SED tasks. Additionally, AVEL data allows models to learn corresponding audio representations for many visual actions in TAL and learn visual cues for sound events in SED, thereby facilitating improvements in both TAL and SED tasks. Besides, some recent task-specific methods [16, 17, 18] have verified that integrating visual and audio modalities is beneficial for both TAL and SED tasks.
@@
In this work, we aim to develop a unified framework to localize all three types of instances in untrimmed videos, solving three temporal video localization tasks (TAL, AVEL and SED) by a single unified model. However, the main obstacles hindering this attempt lie in two aspects. Firstly, different tasks emphasize different video characteristics and modalities. TAL pays more attention to capturing temporal relationships of actions on the visual track, e.g., the visual movement of “playing ten pins” in Fig. 1. SED is dedicated to the fine-grained understanding of audio signals where the events might be invisible, but still contain critical information, such as a narrating person and out-of-screen running cars. In contrast, AVEL assigns equal importance to both auditory and visual cues to locate modality semantic-aligned snippets. For example, in Fig. 1, we can simultaneously see the image and hear the sound of a man speaking from 1.7s to 10.1s. Secondly, the existing datasets for these tasks exhibit distinct properties with significant domain and duration gaps. For example, ActivityNet [19] for TAL task focuses on human daily activities, while UnAV-100 [15] for AVEL task and DESED [20] for SED task contain events from other domains such as animals, nature and tools. Besides, the instance duration of different datasets varies greatly. For instance, over 50% of sound events in DESED span less than one second, while the instances in ActivityNet and UnAV-100 are much longer, with the longest ones lasting over 7 minutes. Therefore, unifying these three distinct tasks is challenging, and just naive multi-task training may lead to a significant decrease in performance.
@@
To tackle the challenges, we introduce UniAV, a Unified Audio-Visual perception framework that enables the model to not only learn and share mutually beneficial knowledge across tasks and modalities but also capture the unique insights inherent to each task. Specifically, we integrate TAL, SED and AVEL tasks into a cohesive framework from the following three aspects. 1) Unified audio-visual encoding. In order to unify diversity between the data from different tasks and obtain generic audio-visual representations, we first employ the audio and visual encoders from the large pre-trained model [21] to uniformly tokenize input videos from all tasks. Then, the obtained embeddings are fed into an audio-visual pyramid transformer network, enabling the model to capture both very short as well as long instances by the thorough cross-modal fusion at multiple temporal scales. 2) Task-specific experts. Due to the substantial divergence of different tasks, learning distinct knowledge for each task is also critical. Thus, we design task-specific expert layers in the transformer blocks to learn task-specific features for each task by adaptively switching to corresponding experts according to the input data. 3) Unified language-aware classifier. Datasets for different tasks pose their own category sets. Instead of using separate task-specific classification heads, we propose a unified language-aware classifier by tokenizing the class vocabularies with task-specific prompts using the pre-trained semantic-aligned text encoder [21]. Benefiting from this new formulation, our model gains the remarkable flexibility to localize rich instance categories across tasks by simply changing task prompts, greatly alleviating the data scarcity of each task, especially for SED task, i.e., only 10 classes in DESED dataset. Furthermore, it enables our model to have an impressive open-vocabulary capability to localize novel instance categories, even across tasks during inference.
@@
With the unified framework, UniAV can learn from diverse task-specific data and handle three video event localization tasks using a single model parameters. Extensive experiments demonstrate that UniAV outperforms its single-task counterparts and the naive multi-task baseline by a large margin, and achieves superior or on-par performances compared to the latest task-specific models. Besides, we find that multi-task joint training can be an effective pre-training step for the single-task models, leading to further performance gains and setting new state-of-the-art results across all three tasks, i.e., ActivityNet 1.3 [19] ($36.2\%$ average mAP) for TAL, DESED [11] ($61.1\%$ average mAP) for SED, and UnAV-100 [15] ($51.7\%$ average mAP) for AVEL.
@@
  • To the best of our knowledge, our UniAV is the first unified framework that solves temporal action localization, sound event detection and audio-visual event localization tasks within a single model, leading to a holistic understanding of video content in real-world scenarios.
@@
  • We propose a unified audio-visual encoding pipeline to address substantial data discrepancies across tasks, and meanwhile design task-specific experts to effectively capture distinct knowledge for each task.
@@
  • We introduce a unified language-aware classifier with semantic-aligned task prompts, allowing the model to flexibly detect various instances across tasks and have an open-set localization capability during inference.
  • Experimental results show that UniAV, with its unified architecture, significantly outperforms both single-task models and the naive multi-task baseline, achieving state-of-the-art performances across all three task benchmarks.
II Related Work
II-A Video Event Localization Tasks
@@
Temporal action localization (TAL) aims to localize and classify action instances occurring in an untrimmed video. Supervised learning-based TAL can be categorized into two-stage and single-stage methods. Two-stage methods [7, 2, 22, 23, 24, 10] first generate action proposals with confidence scores, and then classify their corresponding action segments and refine the temporal boundaries. By contrast, single-stage pipelines [25, 9, 26, 8, 27, 28] simultaneously predict temporal boundaries and action categories for each instance. Besides, some works explore new mechanisms to solve TAL. Yang et al. [29] propose adapters into short-term Vision transformer (ViT) [30] for end-to-end TAL without feature extraction. Foo et al. [31] propose an image-diffusion framework to re-cast TAL as an image generation process. These previous works only focus on temporal modeling within the visual modality (e.g., RGB frames and optical flow), ignoring the information in its corresponding audio track. In recent years, some works [16, 17, 32] attempted to utilize the audio modality in videos for TAL, and found it very helpful to detect the actions that have strong audio cues, thus boosting the model performance.
@@
Sound event detection (SED) is a popular task in the audio signal processing community, which involves temporally detecting sound events in purely acoustic scenes. The DCASE Challenge [33] examines sound event detection in domestic environments as one of the challenge tasks [34, 11, 12]. However, sound events typically come with their corresponding visual information, e.g., the sound events of man speaking, playing guitar, and dog barking. It has been verified that incorporating visual modality is beneficial for SED tasks in some recent works [18, 35].
@@
Audio-visual event localization (AVEL) aims to detect events that are simultaneously audible and visible in video content. Tian et al. [6] introduced the first AVEL dataset and proposed an audio-guided visual attention model for the task. Subsequent studies [13, 36, 37, 38, 4, 14] primarily concentrated on event information modeling and cross-modal fusion strategies, and Zhou et al. [39] extended AVEL to a new open-vocabulary setting. However, these works formulate AVEL as a segment-level classification problem based on trimmed, short video clips. Each video clip only contains a single audio-visual event, which deviates from real-life scenarios involving diverse untrimmed videos. To solve the problem, Geng et al. [15] built the UnAV-100 dataset for localizing audio-visual events in untrimmed videos and proposed a model to recognize multiple events and regress their temporal boundaries in a single pass. Furthermore, few approaches for audio-visual video parsing [40, 41, 42, 43] and multisensory temporal event localization [44] strive to identify audio-only, visual-only and audio-visual events in videos. However, they are all segment-level classification methods with fragmented task definitions, and confined to weakly-supervised settings due to the lack of temporal annotations in videos during training. Besides, the very limited instance categories, i.e., only 25 classes in LLP dataset [40], make the trained model incapable of detecting diverse instances in complex real-life scenarios. In comparison, we develop a multi-task supervised framework that learns multi-modal event localization from large-scale individually collected datasets for TAL, SED and AVEL tasks with rich label vocabularies. Moreover, our model is proposal-based, capable of flexibly regressing and recognizing all temporal instances in untrimmed videos with various lengths.
II-B Multi-Task Learning
@@
Instead of separate training for each individual task, multi-task learning involves tackling multiple related tasks simultaneously, aiming to share and leverage knowledge across them. For example, recent works [45, 46, 47] proposed to jointly learn visual grounding tasks in a collaborative model. Yan et al. [48] presented Unicorn to solve four object tracking problems. Artacho et al. [49] introduced UniPose+ to unify 2D and 3D human pose estimation tasks by a single-pass multi-scale architecture. Lu et al. [50] undertook training across 12 vision-language tasks, with each task having its own task-specific prediction head. For video understanding tasks, UniVTG [51] stands out for its emphasis on unifying three temporal grounding tasks: moment retrieval, highlight detection, and video summarization, employing specific query types. Addressing episodic memory tasks, MINOTAUR [52] exhibits proficiency in handling three egocentric vision tasks [53] through a singular model. Furthermore, recent multimodal large language models (MLLMs) [54, 55, 56, 57, 58, 59] show promise in unified multi-task learning. For instance, VideoLLaMA [55] handles video question answering and captioning tasks simultaneously, and Trace [57] uniformly formats time-aware video tasks into text generation. However, LLM-based methods consume much more computational resources and often underperform compared to non-LLM methods. In this work, we unify temporal action localization, sound event detection, and audio-visual event localization tasks for the first time to achieve diverse modality-aware instance localization in untrimmed videos for holistic video understanding. This is a distinct and pressing problem that was previously overlooked but holds important practical implications in real-world audio-visual scenarios.
III The UniAV Framework
@@
Our goal is to develop a unified framework to localize visual actions, sound events and audio-visual events in an untrimmed video. To achieve this, we propose to unite three video localization tasks: TAL, SED and AVEL, leveraging inherent similarities among them. An overview of the proposed framework is illustrated in Fig. 2.
@@
Figure 2: The overview of our unified framework. Given a visual and audio pair from an untrimmed video, we first tokenize them by a pair of pre-trained visual and audio encoders to obtain generic audio and visual representations. Then, the encoded features are fed into an audio-visual pyramid transformer, which consists of $L_{1}$ uni-modal and $L_{2}$ cross-modal transformer blocks for temporal relation modeling and audio-visual fusion at multiple temporal scales. The task-specific experts inserted in transformer blocks learn distinct knowledge for each task, which is illustrated on the right side. Besides, the categories of each task are encoded with task prompts to compute similarities with the obtained audio-visual feature pyramid, which is used to perform language-aware classification. Finally, the model recognizes classes and regresses temporal boundaries for all types of instances occurring in the video.
Problem statement. Given an untrimmed video containing visual and audio tracks, our model aims to recognize the instance categories and locate the start/end timestamps of all modality-aware instances occurring in the video. Different from the previous fragmented task definitions [8, 6, 34], we uniformly formulate all three tasks (TAL, SED and AVEL) as a joint classification and regression problem.
@@
Formally, we first divide the input video into $T$ visual snippets $\{V_{t}\}_{t=1}^{T}$ and audio snippets $\{A_{t}\}_{t=1}^{T}$, where $T$ varies across videos. For the given task, the groundtruth of an input video is denoted as $Y=\{y_{n}=(t_{s,n},t_{e,n},c_{n})\}^{N}_{n=1}$, where $t_{s,n}$, $t_{e,n}$ are the start and end timestamps of $n$-th instance, $c_{n}\in\{1,\cdots,C\}$ is the instance category, and $N$ is the number of instances occurring in the video. Following the setting [8, 15], the model is expected to predict $\hat{Y}=\{\hat{y}_{t}=(d_{s,t},d_{e,t},p(c_{t}))\}_{t=1}^{T}$ for each moment during inference, where $p(c_{t})\in\mathbb{R}^{1\times C}$ is the probabilities of $C$ categories at moment $t$, $d_{s,t}$ and $d_{e,t}$ are the distances between the moment $t$ to the instance’s start and end timestamp, respectively. $d_{s,t}$ and $d_{e,t}$ are only defined when an instance presents at moment $t$. Therefore, the prediction results for the given task can be decoded from $\hat{y}_{t}=(d_{s,t},d_{e,t},p(c_{t}))$ by:
$c_{t}={\arg\max}\,p(c_{t}),\quad t_{s,t}=t-d_{s,t},\quad t_{e,t}=t+d_{e,t}.$ (1)
III-A Unified Audio-Visual Encoding
Audio and visual signals are equally crucial for AVEL [6, 15], and both serve as important complementary information for TAL and SED to boost the localization performances [16, 35, 18]. Therefore, to solve these three tasks together using a single unified framework, we propose a unified audio-visual encoding pipeline to utilize both audio and visual modalities of all input videos for all tasks. The proposed pipeline consists of two encoding processes as follows: